OpenStudy (anonymous):

The number of Math homework questions is seven more than one−half the number of History homework questions. There are 15 Math homework questions. Which of the following equations would you use to find the number of History homework questions? Note: h represents the number of questions on your History homework.

OpenStudy (anonymous):

15/2+7=h 15+7/2=h h+7/2=15 1/2h+7=5 These are your choices..

OpenStudy (anonymous):

1/2h+7=15? Are you sure the last one isnt 15?

OpenStudy (anonymous):

You had it right.. i miss typed the last option..

OpenStudy (anonymous):

Awesome.

OpenStudy (anonymous):

What class on FLVS are you taking?